OpenStudy (anonymous):

Solve the following system of equations: 2x + 3y − z = 1 3x + y + 2z = 12 x + 2y − 3z = −5

OpenStudy (imstuck):

Which method are you using in class? Are you doing matrices? Or elimination of a variable?

OpenStudy (anonymous):

I can use elimination or substitution.

OpenStudy (imstuck):

Ok, well elimination is a bunch easier! Can we do that?

OpenStudy (anonymous):

yes

OpenStudy (imstuck):

The goal is to take the first 2 equations and eliminate a variable, then take the second and third equation and eliminate the same variable, then the two resulting equations, each of them minus one variable can be solved simultaneously for x or y or z. Then use that to find the other variables. Let's start with eliminating the z. Cuz z's are weird in equation, ok? ; )

OpenStudy (imstuck):

We will eliminate the z between the first two equations: 2x + 3y - z = 1 3x + y + 2z = 12

OpenStudy (imstuck):

We have to make the coefficients the same number, but one of them negative and the other positive so they cancel each other out. In order to do that, let's make both the coefficients in front of those z's a 2.

OpenStudy (imstuck):

To do that, multiply the first equation by a 2 like this:

OpenStudy (imstuck):

2(2x + 3y - z = 1) --> 4x + 6y - 2z = 2

OpenStudy (imstuck):

Now let's add that to the second equation and see what happens:

OpenStudy (imstuck):

4x + 6y - 2z = 2 3x + y + 2z = 12 ---------------- 7x + 7y + 0z = 14 Of course 0z is equal to no z's at all, so what we are left with is: 7x + 7y = 14

OpenStudy (imstuck):

That's the first 2 equations. Now let's do equation 2 and 3 the same way:

OpenStudy (imstuck):

3x + y + 2z = 12 x + 2y - 3z = -5

OpenStudy (imstuck):

In order to eliminate the z's in these two, we have to again get the coefficients the same, but one negative and one positive. Between a 2 and a 3, 6 is the LCM, so we have to multiply the second equation by a 3, and the third equation by a 2 to make them both 6's. Like this:

OpenStudy (imstuck):

3(3x + y + 2z = 12) 2( x + 2y - 3z = -5)

OpenStudy (imstuck):

Multiplying those through you get this:

OpenStudy (imstuck):

9x + 3y + 6z = 36 2x + 4y - 6z = -10 ----------------- 11x + 7y + 0z = 26 Of course, again, the 0z is gone, so we have two new equations now, which are:

OpenStudy (imstuck):

7x + 7y = 14 11x + 7y = 26

OpenStudy (imstuck):

Now we have to work to eliminate another variable, which is easy this time...there are 7's in front of the both y coordinates, but both of them are positive. We have to make one of them negative, so we will multiply either one by a -1, like this (I am going to pick the top one to multiply by a -1):

OpenStudy (imstuck):

-1(7x + 7y = 14) --> -7x - 7y = -14

OpenStudy (imstuck):

Adding that to the bottom equation we now have this:

OpenStudy (imstuck):

-7x - 7y = -14 11x + 7y = 26 ------------- 4x + 0y = 12 or 4x = 12.

OpenStudy (imstuck):

Solve that for x and get that x = 3. NOW fill the 3 back in to one of our equations that has only an x and a y in it to get this (I m going to use the equation 7x + 7y = 14):

OpenStudy (imstuck):

7(3) + 7y = 14 --> 21 + 7y = 14 --> 7y = 14 - 21 --> 7y = -7 --> y = -1.

OpenStudy (imstuck):

Now we have x = 3 and y = -1. Fill those into one our original equations with all three variables to solve for z:

OpenStudy (imstuck):

2x + 3y - z = 1 --> 2(3) + 3(-1) - z = 1 --> 6 - 3 - z = 1 --> 3 - z = 1 --> -z = 1 - 3 --> -z = -2 or z = 2.

OpenStudy (imstuck):

Now you have all your variables: x = 3, y = -1, z = 2. See that?

OpenStudy (anonymous):

yes! thank you!

OpenStudy (imstuck):

You are welcome!